When an empty lift is moving down with an acceleration of g/4ms-2, the tension in the cable is 9000N. When the lift is moving up with an acceleration of g/3ms-2, the tension in the cable is,

Grade:11

2 Answers

Kshitiz goel
6 years ago
When elevator is going down ....Acc of elevator downward is g/4.Tension on the elevator is upward T=9000N Also downward g acting on the elevator gNow by carefully balancing these forces we get T= Mg + Mg/4 Where M = mass of elevator....From here we get mass of elevatir as 720 kg ..Now case 2On elevator Tension is upward ...Acceleration g/3 upwardsAlso downward gravitaional acceleration g Now carefully balncing the forces we get,T + Mg/3 = MgT= 4800NI suppose the answer is 4800N...Thank u
Vikas TU
6 years ago
When the lift is going down, the increasing speed is g/4m/s2. 
Descending pressure = 9000N, 
At the point when the lift is going up, increasing speed = g/4 m/s2 
At that point now adjusting the Force, 
T = Mg+ Mg/4 
    M = mass of lift. 
    g = 10 m/s2 
9000 = 10M +2.5M 
9000 = 12.5 M 
    M = 90000/125 
    M = 720 kg 
At the point when the lift is going down, quickening is g/3 m/s2 
Presently the compel is adjusted, 
T + Mg/3 = Mg 
T + (7200)/3 = 7200 
T+2400 = 7200 
T = 4800 N